OpenStudy (anonymous):

How many hours does it take to drive from Trenton, NJ to Harrisburg at 55 mph? Round to the nearest tenth hour.

OpenStudy (anonymous):

approx. distance=150 m speed=55m/h time=d/s=150/55=?